Probabilities and Counting. Yahtzee is a game that involves six fair dice. When rolling all six...

  1. Probabilities and Counting. Yahtzee is a game that involves six fair dice. When rolling all six dice, what is the probability that at least two of them will show the same number?
Answer #1

P(at least two of them will show the same number) =1-P(all different)

= 1- (6!/(6^6))

=319/324 = 0.984568